2.00 grams of magnesium carbonate are reacted with excess hydrochloric acid. If the density of carbon dioxide gas at atmospheric conditions is 24 L/mol, what is the volume of the gas evolved?
1
Expert's answer
2020-10-26T14:49:41-0400

The reaction equation:

MgCO3 + 2HCl = MgCl2 + CO2 + H2O

The quantity of magnesium carbonate:

2.00 g / (84.31 g/mol) = 0.0237 mol

The quantity carbon dioxide is the same (according to the balanced equation above).

The volume of carbon dioxide:

0.0237 mol * 24 L/mol = 0.57 L

Answer: 0.57 L


Please note! that the "density" of any gas at atmospheric conditions is 22.4 L/mol (not 24!).

In this case the answer is: 0.0237 mol * 22.4 L/mol = 0.53 L